# 3QFP: Efficient neural implicit surface reconstruction using Tri-Quadtrees and Fourier feature Positional encoding

### 13 Jan 2024 v1

@@ Abstract
Neural implicit surface representations are curre ntly receiving a lot of interest as a means to achieve high-fidelity surface reconstruction at a low memory cost, compared to traditional explicit representations. However, state-of-the-art methods still struggle with excessive memory usage and non-smooth surfaces. This is particularly problematic in large-scale applications with sparse inputs, as is common in robotics use cases. To address these issues, we first introduce a sparse structure, tri-quadtrees, which represents the environment using learnable features stored in three planar quadtree projections. Secondly, we concatenate the learnable features with a Fourier feature positional encoding. The combined features are then decoded into signed distance values through a small multi-layer perceptron. We demonstrate that this approach facilitates smoother reconstruction with a higher completion ratio with fewer holes. Compared to two recent baselines, one implicit and one explicit, our approach requires only 10%–50% as much memory, while achieving competitive quality.
@@ III Method
Our method learns a continuous signed distance function representation of the environment given lidar scans and known poses. Specifically, the world coordinate $\bm{p}_{i}\in\mathbb{R}^{3}$ is mapped into a SDF value $s_{i}\in\mathbb{R}$. As shown in fig. 2, our neural implicit representation is composed of two components: the learnable features stored in the quadtree nodes and a globally shared MLP for predicting the SDF value. The features and the network parameters are optimized in real-time by using direct lidar measurement to supervise network predictions.
Figure 2: Overview of our method. We represent the scene with three planar quadtrees $\mathcal{M}_{i}^{\ell}$, $i\in\{XZ,YZ,XY\}$ and $\ell$ represents the quadtree depth. We store features in the deepest $H$ levels of resolution of quadtrees. When querying for a point $\bm{p}$, we project it onto planar quadtrees to identify the node containing $\bm{p}$ at the level $\ell$. The feature of $\bm{p}$ is then calculated by bilinear interpolation based on the queried location and vertex features. We add features at the same level and concatenate among different levels. Concatenated with the positional encoding $\gamma(\bm{p})$, $\bm{p}$’s feature ($\Phi(\bm{p})$) is fed into a small MLP ($\mathcal{F}_{\Theta}$) to predict the SDF value. The learnable features stored in the quadtree nodes and the network parameters are optimized in real-time using the loss function $\mathcal{L}_{\text{bce}}$. The learnable feature vectors have length $d$ and the positional encoding feature vector has length $6m$.
@@ III-A Tri-Quadtrees feature representation
Tri-Quadtrees: While storing learnable features in spatial voxel grids [34] can lead to fast convergence rates, dense 3D grids suffer from a cubical growth rate of memory usage as the environment size increases. To avoid storing unnecessary features in free space, prior work [20, 23, 16, 24] employ octree to store features only within voxel grids where surface points are located. Other research such as [21] adopts feature planes, which represent scene features by three orthogonal planes. In this paper, we propose a novel data structure to represent spatial features—tri-quadtrees, which is more compact than previous methods.
Specifically, given lidar scans and known poses, we first project the 3D point clouds onto three axis-aligned orthogonal 2D planes and then construct quadtrees on each plane with a maximum depth of $L_{\text{max}}$ and resolution of the leaf nodes equal to $r$. Unlike the method in NGLoD [16], which utilises all levels of resolution in the octree to store features, we only use the deepest $H$ levels of quadtrees for feature representation to balance the quality of the reconstruction and memory footprint. That is, only levels $\ell\in\{L_{\text{max}}-H+1,\cdots,L_{\text{max}}\}$ in the quadtree are used to interpolate features. By default, we set $H=3$, which was empirically sufficient shown to achieve good results.
@@ III-B Fourier features positional encoding
Although the learnable features can contribute to a more accurate reconstruction, they do not provide the necessary hole filling and smoothness, as demonstrated by [26]. To improve the completion ratio of the reconstruction, we combine the learnable features and positional encoding. We demonstrate through experiments (section IV-D) that combining learnable features with the positional encoding can boost the completion ratio with only a marginal increase in computational cost.
@@ IV-A Experiment Setup
Implementation Details: In all experiments, our method uses the parameters shown in table I. We set $\sigma^{2}=50$ by parameter searching on MaiCity and apply it on both datasets.

### 07 Apr 2024 v2

@@ Abstract
Neural implicit surface representations are currently receiving a lot of interest as a means to achieve high-fidelity surface reconstruction at a low memory cost, compared to traditional explicit representations. However, state-of-the-art methods still struggle with excessive memory usage and non-smooth surfaces. This is particularly problematic in large-scale applications with sparse inputs, as is common in robotics use cases. To address these issues, we first introduce a sparse structure, tri-quadtrees, which represents the environment using learnable features stored in three planar quadtree projections. Secondly, we concatenate the learnable features with a Fourier feature positional encoding. The combined features are then decoded into signed distance values through a small multi-layer perceptron. We demonstrate that this approach facilitates smoother reconstruction with a higher completion ratio with fewer holes. Compared to two recent baselines, one implicit and one explicit, our approach requires only 10%–50% as much memory, while achieving competitive quality. The code is released on https://github.com/ljjTYJR/3QFP.
@@ III Method
Our method learns a continuous signed distance function representation of the environment given lidar scans and known poses. Specifically, the world coordinate $\bm{p}_{i}\in\mathbb{R}^{3}$ is mapped into a SDF value $s_{i}\in\mathbb{R}$. As shown in fig. 2, our neural implicit representation is composed of two components: the learnable features stored in the quadtree nodes and a globally shared MLP for predicting the SDF value. The features and the network parameters are learned during test time by using direct lidar measurement to supervise network predictions.
Figure 2: Overview of our method. We represent the scene with three planar quadtrees $\mathcal{M}_{i}^{\ell}$, $i\in\{XZ,YZ,XY\}$ and $\ell$ represents the quadtree depth. We store features in the deepest $H$ levels of resolution of quadtrees. When querying for a point $\bm{p}$, we project it onto planar quadtrees to identify the node containing $\bm{p}$ at the level $\ell$. The feature of $\bm{p}$ is then calculated by bilinear interpolation based on the queried location and vertex features. We add features at the same level and concatenate among different levels. Concatenated with the positional encoding $\gamma(\bm{p})$, $\bm{p}$’s feature ($\Phi(\bm{p})$) is fed into a small MLP ($\mathcal{F}_{\Theta}$) to predict the SDF value. The learnable features stored in the quadtree nodes and the network parameters are learned by test-time optimization using the loss function $\mathcal{L}_{\text{bce}}$. The learnable feature vectors have length $d$ and the positional encoding feature vector has length $6m$.
@@ III-A Tri-Quadtrees feature representation
Tri-Quadtrees: While storing learnable features in spatial voxel grids [34] can lead to fast convergence rates, dense 3D grids suffer from a cubical growth rate of memory usage as the environment size increases. To avoid storing unnecessary features in free space, prior work [20, 23, 16, 24] employs octree to store features only within voxel grids where surface points are located. Other research such as [21] adopts feature planes, which represent scene features by three orthogonal planes. In this paper, we propose a novel data structure to represent spatial features—tri-quadtrees, which is more compact than previous methods.
Specifically, given lidar scans and known poses, we first project the 3D point clouds onto three axis-aligned orthogonal 2D planes and then construct quadtrees on each plane with a maximum depth of $L_{\text{max}}$ and resolution of the leaf nodes equal to $r$. Unlike the method in NGLoD [16], which uses all levels of resolution in the octree to store features, we only use the deepest $H$ levels of quadtrees for feature representation to balance the quality of the reconstruction and memory footprint. That is, only the levels $\ell\in\{L_{\text{max}}-H+1,\cdots,L_{\text{max}}\}$ in the quadtree are used to interpolate features. By default, we set $H=3$, which was empirically sufficient to achieve good results.
@@ III-B Fourier features positional encoding
Although the learnable features can contribute to a more accurate reconstruction, they do not provide the necessary hole filling and smoothness, as demonstrated by [26]. To improve the completion ratio of the reconstruction, we combine the learnable features and the positional encoding. We demonstrate through experiments (section IV-D) that combining learnable features with the positional encoding can boost the completion ratio with only a marginal increase in computational cost.
@@ IV-A Experiment Setup
Implementation Details: In all experiments, our method uses the parameters shown in table I (we use these parameters to trade off efficiency and accuracy). We set $\sigma^{2}=50$ by parameter searching on MaiCity and apply it on both datasets.